# Exploring Representations and Interventions in Time Series Foundation Models

### v2

@@ 1 Introduction
Foundation models have taken significant strides in modeling both textual (Brown et al., 2020) and visual (Dosovitskiy et al., 2020) data, and have made complex language and image processing accessible to non-experts. These models are pre-trained on massive internet-scale datasets and can be used to solve multiple tasks across a variety of domains, with little to no adaptation. Recently, a growing body of work (Garza and Mergenthaler-Canseco, 2023; Goswami et al., 2024; Rasul et al., 2024; Das et al., 2024; Woo et al., 2024; Ansari et al., 2024) has extended the benefits of this paradigm to time series data, a modality prevalent in fields such as finance (Taylor, 2008), healthcare (Goswami et al., 2021), and climate science (Schneider and Dickinson, 1974).
@@
Time series foundation models (TSFMs) have shown promising performance on multiple modeling tasks such as forecasting, classification, anomaly detection, and imputation, across a wide range of domains, and in settings with varying amounts of data and supervision. However, the underlying mechanisms and learned representations of TSFMs remain largely unexplored. Little is known about the characteristics of learned representations, the kinds of concepts that these models are learning, and how can these concepts could be manipulated to influence model outputs. A deeper understanding of the inner workings of TSFMs is key to enhancing their performance and trustworthiness.
We address these knowledge gaps, by systematically probing TSFMs and intervening in them. First, we examine these models through the lens of representational similarity, identify learned concepts, and localize them to specific hidden states. We find that TSFMs learn redundant representations, and intuitive concepts such as base patterns (e.g., constant vs. sinusoidal waves), amplitudes, periodicity, and trends; often in specific layers and patches. We leverage these insights to intervene in TSFMs in two ways: (1) we exploit the redundant representations to prune models, reducing their size and improving inference speed while preserving accuracy, and (2) we steer model predictions along conceptual directions (e.g., adding an upward trend to forecasts) using synthetic time series. Our work is the first step towards understanding the inner workings of TSFMs and leveraging this knowledge to develop strategies for improving their capabilities.
@@ 2 Related Work
Time Series Foundation Models (TSFMs). TSFMs are versatile neural networks pre-trained on vast amounts of time series data, and have shown remarkable capabilities in producing accurate predictions even in zero-shot settings. Recently, a number of TSFMs have been proposed (Garza and Mergenthaler-Canseco, 2023; Liu et al., 2023; Das et al., 2024; Woo et al., 2024; Goswami et al., 2024; Ansari et al., 2024; Ekambaram et al., 2024; Talukder et al., 2024). While most TSFMs are based on variations of the Transformer architecture (Vaswani, 2017), they exhibit notable differences in tokenization strategies, pre-training datasets, and the specific tasks they are designed to address. While our methods are broadly applicable to Transformer-based TSFMs, we will primarily focus on analyzing MOMENT Goswami et al. (2024), Chronos Ansari et al. (2024), and MOIRAI Woo et al. (2024), three representative TSFMs that are fully open-source and offer distinct approaches to time series tokenization (patch vs. discrete), architecture (encoder-only vs. encoder-decoder), and pre-training objectives (imputation vs. forecasting).
@@
Analyzing Representations of Deep Learning Models. Deep learning models often operate as black boxes, making their internal mechanisms and learned representations difficult to understand. One approach to gaining insights into these models is by comparing their intermediate representations. Similarity metrics can be used to determine the similarity or dissimilarity of representations at different stages of a model, revealing the hierarchy, homogeneity, and redundancy of learned features. Previous studies focused on quantifying the similarity of neural network representations Raghu et al. (2017); Kornblith et al. (2019). Raghu et al. (2021) demonstrated the use of these metrics for analyzing and comparing representations in vision transformers and CNNs, providing valuable insights into their functioning. Nguyen et al. (2021) investigated the impact of model size and training data ratios on similarity patterns by varying the depth and width of different models. They also explored model pruning based on similarity. Building on these these studies, our work presents the first comprehensive analysis of representations learned by TSFMs, offering useful insights into their internal workings.
@@
Identifying and Manipulating Learned Concepts in Pre-trained Models. Understanding the internal representations learned by pre-trained models has been an active area of research, particularly in the context of LLMs and vision models. Previous studies have explored whether individual neurons or directions in a model’s latent space correspond to specific features or concepts (Dalvi et al., 2019; Goh et al., 2021; Gurnee et al., 2023; Elhage et al., 2022). These investigations often focus on identifying linear representations, where features are encoded as linear combinations of neuron activations. Recent work has also employed various probing techniques to classify and interpret these internal representations, addressing aspects such as truthfulness and model robustness (Azaria and Mitchell, 2023; Zou et al., 2023; Burns et al., 2023; Marks and Tegmark, 2023). While many of these studies rely on meticulously curated datasets to probe language and vision models, we demonstrate that for time series models, synthetic data generated using simple mechanisms can effectively identify, localize, and probe the concepts learned by these models.
3 Probing and Intervening in Time Series Foundation Models
@@
We study TSFMs using two complementary analysis and intervention methods. In Section 3.1, we examine learned representations through the lens of similarity, uncovering the redundancy inherent in TSFM representations. We leverage this redundancy to prune multiple layers of pre-trained models, thereby improving their efficiency without compromising accuracy. In Section 3.2, we identify the specific concepts learned by TSFMs, localizing them to specific hidden states. Furthermore, we explore the ability to steer model predictions along these conceptual directions, enabling us to influence model outputs in targeted ways.
3.1 Analyzing Representational Similarity for Effective Pruning
@@
To gain a comprehensive understanding of the similarity between learned representations, we considered several metrics commonly employed in the literature. While our primary analysis relies on Centered Kernel Alignment (CKA) Kornblith et al. (2019), we also explored Cosine Similarity and Singular Vector Canonical Correlation Analysis (SVCCA) (Raghu et al., 2017). For brevity, we provide a brief overview of CKA below, while detailed descriptions of the remaining metrics can be found in Appendix B.
Representational Similarity using Centered Kernel Alignment (CKA).
@@
where HSIC denotes the Hilbert-Schmidt Independence Criterion (Gretton et al. (2005)).
Figure 1: For each identified block of layers exhibiting redundant representations (red), we remove the internal layers of the block by zeroing out their weights (blue). For example, if a block consists of five layers, we prune layers 2 through 4, retaining only the first and last layers to reduce representation redundancy while maintaining model integrity.
@@
1:  for each block $b_{i}$ in $\mathcal{B}$ do
2:     Let $b_{i}$ consist of layers $l_{s}$ to $l_{e}$Block edges at $l_{s}$ and $l_{e}$}
3:     for layer index $j=s+1$ to $e-1$ do
4:        Zero out the weights of layer $l_{j}$ in model $\mathcal{M}$
5:     end for
6:  end for
7:  return  Pruned model $\mathcal{M}^{\prime}$
Pruning TSFMs Based on Representational Similarity.
Large TSFMs typically learn redundant representations, which often manifest as block-like structures in heatmaps depicting pairwise similarity between layer activations (Figure 1). We can leverage this redundancy to downsize TSFMs, to improve their inference speed, without affecting their accuracy. We build on prior work (Nguyen et al., 2021) and propose a simple layer pruning strategy, which we call Block-wise Pruning, outlined in Algorithm 1. To preserve the structural integrity of each block, we retain the first and last layers of each block while zeroing out the weights of the intermediate layers. The skip connections within transformer blocks ensure that signals and gradients continue to flow through the rest of the network.
Research Questions and Experimental Setup.
To gain a deeper understanding of TSFM representations, we investigate the following research questions: (RQ1) How similar are the representations learned by different layers within a single model? (RQ2) How do these representations vary across models of different sizes? (RQ3) How similar are representations learned by different TSFMs from the same family? To answer these questions, we use Centered Kernel Alignment (CKA) to measure the similarity between representations at different layers of TSFMs, and visualize the results using heatmaps.
@@
To demonstrate the effectiveness of our proposed pruning strategy, we explore two pruning configurations, one in which prune all redundant blocks, and the other where prune only a single block. We compare the performance of these pruned models to the original, unpruned TSFMs using standard task-specific accuracy metrics (Mean Squared Error and Mean Absolute Error) and efficiency metrics (inference time in milliseconds and theoretical model size in megabytes). We evaluate these models on widely used imputation (Zhou et al., 2021) and forecasting (Ansari et al., 2024) benchmarks in both zero-shot settings and after linear probing (Goswami et al., 2024).
@@
Figure 2: Overview of linear probing, concept localization, and steering. Linear probing involves training separate linear models for each layer or layer and patch to classify time series $\mathbf{x}$ into constant $c$ and sinusoid $s$ classes. Classifiers $f_{ij}(\mathbf{h}_{i}^{(j)},\theta_{i}^{(j)})$ are trained on the hidden representation $\mathbf{h}_{i}^{(j)}$ at each $i$-th layer and $j$-th token to update the parameters $\theta_{i}^{j}$. Concept localization is achieved by computing Fisher’s Linear Discriminant Ratio (LDR) between the classes at each layer and token using mean and variance statistics of $h_{i}^{(j)}$ for each predicted class, $\hat{y}$. The LDR output is scaled between 0 and 1 using min-max scaling to allow for consistent comparison across layers. Concept steering vector can be derived for each $i$-th layer by calculating the difference between the median activation matrices of the sinusoid and constant time series classes, $\mathbf{M}_{i_{s}}-\mathbf{M}_{i_{c}}$ and then stacked into a steering matrix $\mathbf{S}$ for the whole model. During model inference, the steering matrix can be used to steer model predictions towards desired concepts, or classes, by updating the embeddings as $\mathbf{h}_{i}\leftarrow\mathbf{h}_{i}+\alpha\mathbf{S}_{i}$, where $\alpha$ is a scalar that controls the strength of the intervention.
3.2 Identifying, Localizing and Steering Time Series Concepts
Generating Synthetic Data for Conceptual Analysis of TSFMs.
To systematically explore the ability of TSFMs to understand intuitive time series concepts, we randomly generate a large number of synthetic univariate time series. Each randomly generated time series belong to one of two pattern classes: constant or sinusoidal. Constant patterns, represented by $y(t)=\alpha t+\beta$, capture long-term non-periodic trends. Sinusoidal patterns, modeled as $y(t)=A\sin\left(\frac{2\pi t}{P}\right)$, represent periodic processes. By controlling the parameters $\alpha$, $\beta$, $A$, and $P$, we can systematically generate time series with varying slope, intercept, amplitude, and periodicity. Despite their simplicity, these data generation mechanisms capture a wide range of real-world time series patterns. For a detailed description of the data generation process, please refer to Appendix A.
In this section, we build on the investigation approach outlined in (Marks and Tegmark, 2023). We say that a feature is linearly represented in a foundation model $\mathcal{M}$ if it is represented as a direction in its latent space. As a concrete example, consider that we want to identify whether $\mathcal{M}$ can distinguish between constant and sinusoidal patterns. If this feature is linearly represented in $\mathcal{M}$, we also want to identify which layer $l$ in $\mathcal{M}$ learns this concept in the most discriminant way.
(a) Constant with increasing trend
(b) Sinusoid signal
(c) Sinusoid with increasing trend
@@ 3.2 Identifying, Localizing and Steering Time Series Concepts
Figure 3: Examples of synthetic data generated for experiments include constant signals with varying trend (a), sinusoidal signals with varying frequency (b), and compositions of constant signals with varying trends and sinusoidal signals, resulting in sinusoidal signals with varying trends (c). Synthetic data is used in linear probing, concept localization, and concept steering experiments.
@@
To determine whether the feature (sinusoidal vs. constant time series) is linearly represented, we first generate a synthetic time series dataset by varying this feature. This dataset comprises of multiple sinusoids and constant time series randomly sampled using our data generating function. Using this dataset, we extract intermediate representations of each time series from the residual stream of each layer. Let $\mathbf{h}_{i}^{(j)}\in\mathbb{R}^{n\times D}$ denote the hidden representation of a time series $\mathbf{x}$ at $i$-th layer and $j$-th token of $\mathcal{M}$, where $D$ is the dimensionality of the hidden layer. Linear probing involves training separate linear models for each layer and token to classify time series $\mathbf{x}$ as a constant or sinusoid pattern. Classifiers $f_{ij}(\mathbf{h}_{i}^{(j)},\theta_{i}^{(j)})$ are trained on the hidden representation $\mathbf{h}_{i}^{(j)}$ at each $i$-th layer and each $j$-th token to update the parameters $\theta_{i}^{j}$. Additionally, we perform probing on representations averaged along the token dimension for each $i$-th layer. The linear probes are trained to optimize the Fisher Criterion, a function that aims to maximize the distance between class means while minimizing within-class variance:
@@
Here, $\mu_{s}$ and $\mu_{c}$ correspond to the mean computed across the $n$ dimension for each class. Similarly, $\sigma^{2}_{s}$ and $\sigma^{2}_{c}$ correspond to the variance computed across the sample dimension $n$ for each class. Let $\mathbf{V}=[v_{i,j}]\in\mathbb{R}^{L\times N}$ be the matrix of LDR values, where $v_{i,j}$ represents the LDR value for the $i$-th layer and $j$-th token, with $l$ layers and $N$ tokens. The LDR output is scaled between 0 and 1 using min-max scaling to allow for consistent comparison across layers. By visualizing the scaled LDA values as shown in Figure 2, one can identify which layers and tokens exhibit the highest degree of separation between classes, offering insights into the network’s internal representations for concept intervention techniques.

### v3

@@ 1 Introduction
Foundation models have taken significant strides in modeling both textual (Brown et al., 2020) and visual (Dosovitskiy et al., 2020) data, and have made complex language and image processing accessible to non-experts. These models are pre-trained on massive internet-scale datasets and can be used to solve multiple tasks across a variety of domains, with little to no adaptation. Recently, a growing body of work (Garza & Mergenthaler-Canseco, 2023; Goswami et al., 2024; Rasul et al., 2024; Das et al., 2024; Woo et al., 2024; Ansari et al., 2024) has extended the benefits of this paradigm to time series data, a modality prevalent in fields such as finance (Taylor, 2008), healthcare (Goswami et al., 2021), and climate science (Schneider & Dickinson, 1974).
@@
Time series foundation models (TSFMs) have shown promising performance on multiple modeling tasks such as forecasting, classification, anomaly detection, and imputation, across a wide range of domains, and in settings with varying amounts of data and supervision. However, the underlying mechanisms and learned representations of TSFMs remain largely unexplored. The nature of learned representations in TSFMs remains largely unknown, including (1) their organizational structure, (2) the human-interpretable concepts they encode and their locations, and (3) how they can be manipulated to integrate prior knowledge for more informed predictions. Gaining deeper insight into the inner workings of TSFMs is key to improving their performance and trustworthiness.
We address each of these knowledge gaps by proposing approaches to systematically analyze, probe, and intervene in TSFMs through their learned representations. To address the first gap, we explore the fundamental question: Do different TSFMs learn and organize knowledge in a similar manner? This analysis is crucial for understanding how knowledge is organized in different models and whether representations are redundant across groups of layers, which can be potentially leveraged to improve inference efficiency. Next, we address the second gap by investigating: What human-interpretable concepts can TSFMs learn, and where do these concepts emerge within the latent representations of the model? Building on our layer-wise representation analysis, we propose new methods to identify and localize these concepts across layers. Finally, we tackle the third gap by asking: Can we intervene in TSFMs during inference to guide their output? Through a novel approach, we demonstrate how to steer learned concepts by adjusting representations across layers, allowing the model to generate concept-informed predictions.
The main contributions of this paper are:
  1. (i)
    TSFM Representation Similarity Analysis and Pruning. We translate prior work on vision transformers and convolutional neural networks to quantify the representation similarity in TSFMs. Our analysis reveals that model size substantially influences how representations are organized. In larger models, we observe high CKA similarity across groups of layers, indicating redundant knowledge storage. Leveraging this insight, we propose a novel block-wise layer pruning strategy to remove redundant representations. This approach reduces model size and speeds up inference while maintaining accuracy.
  2. (ii)
    Concept Identification in TSFM Representations. We introduce a new method to identify and localize specific time series concepts, such as trends and seasonal patterns, across model layers. Our findings show that TSFMs gradually learn these intuitive concepts across layers.
  3. (iii)
    Steering Methods for Concept-Informed TSFM Predictions. We demonstrate a novel approach to steer model predictions by intervening in TSFM latent representations along specific conceptual directions (e.g., introducing an upward trend in forecasts) as a way to guide model output toward concept-informed predictions without additional training or fine-tuning.
@@ 2 Related Work
Time Series Foundation Models (TSFMs). TSFMs are versatile neural networks pre-trained on vast amounts of time series data, and have shown impressive accuracy even in zero-shot settings. Several TSFMs have been proposed (Garza & Mergenthaler-Canseco, 2023; Liu et al., 2023; Das et al., 2024; Woo et al., 2024; Goswami et al., 2024; Ansari et al., 2024; Ekambaram et al., 2024; Talukder et al., 2024). Most of these are based on variations of the Transformer architecture (Vaswani, 2017), with differences in tokenization strategies, pre-training datasets, and the specific tasks they are designed to address. Consequently, we focus on developing methods to understand Transformer-based models. We specifically analyze three representative models: MOMENT (Goswami et al., 2024), Chronos (Ansari et al., 2024), and MOIRAI (Woo et al., 2024), which are fully open-source and offer distinct approaches to time series tokenization (patch vs. discrete), architecture (encoder-only vs. encoder-decoder), and pre-training objectives (imputation vs. forecasting). Recent advancements in TSFMs have focused mostly on improving their capabilities through scaling, specialization, or improving inference efficiency (Liu et al., 2024a, 2023), as well as by incorporating long and multivariate contexts (Liu et al., 2024b; Żukowska et al., 2024). However, there has been limited research on understanding what a specific TSFMs learn. For example, Goswami et al. (2024) showed that MOMENT learns some human interpretable concepts, while Ansari et al. (2024) investigated the failure modes of Chronos in controlled settings, and Potosnak et al. (2024) explored compositional reasoning within time series models, including TSFM.
@@
Analyzing Representations of Deep Learning Models. Deep learning models often operate as black boxes, making their internal mechanisms and learned representations difficult to understand. One approach to gaining insights into these models is by comparing their intermediate representations. Similarity metrics can be used to determine the similarity or dissimilarity of representations at different stages of a model, revealing the hierarchy, homogeneity, and redundancy of learned features. Previous studies focused on quantifying the similarity of neural network representations (Raghu et al., 2017; Kornblith et al., 2019). (Raghu et al., 2021) demonstrated the use of these metrics for analyzing and comparing representations in vision transformers and CNNs, providing valuable insights into their functioning. (Nguyen et al., 2021) investigated the impact of model size and training data ratios on similarity patterns by varying the depth and width of different models. They also explored model pruning based on similarity. Building on these these studies, our work presents the first comprehensive analysis of representations learned by TSFMs, offering useful insights into their internal workings.
@@
Identifying Learned Concepts in Pre-trained Models. Understanding the internal representations learned by pre-trained models has been an active area of research, particularly in the context of LLMs and vision models. Previous studies have explored whether individual neurons or directions in a model’s latent space correspond to specific features or concepts (Dalvi et al., 2019; Goh et al., 2021; Gurnee et al., 2023; Elhage et al., 2022). These investigations often focus on identifying linear representations, where features are encoded as linear combinations of neuron activations. Recent work has also employed various probing techniques to classify and interpret these internal representations, addressing aspects such as truthfulness and model robustness (Azaria & Mitchell, 2023; Zou et al., 2023; Burns et al., 2023; Marks & Tegmark, 2023). While prior work has focused on language and vision models, we show that with simple techniques, we can effectively probe TSFMs to identify and localize concepts.
3 Methods
@@
We study TSFMs using three complementary methods of analysis and interventions concerning model layer-wise representations. In Section 3.1, we examine learned representations through the lens of similarity, uncovering stored knowledge redundancy in TSFM representations. We leverage this redundancy to prune multiple layers of pre-trained models, thereby improving their efficiency. In Section 3.2, we identify specific concepts learned by TSFMs and localize them to specific parts of model layers and tokens. In Section 3.3, we explore the ability to steer model predictions by intervening in model hidden representations across layers, enabling us to guide model output toward concept-informed predictions.
3.1 Representation Similarity Analysis and Pruning
To gain a deeper understanding of TSFM representations, we investigate the following research questions: (RQ1) How similar are the representations learned by models of the same size but belonging to different families? (RQ2) How do these representations differ across models of varying sizes within the same family? (RQ3) How similar are the representations learned by corresponding layers of different TSFMs within the same family?
@@ 3.1 Representation Similarity Analysis and Pruning
We consider several metrics commonly employed in the literature to analyze the similarity between learned representations. While our primary analysis relies on Centered Kernel Alignment (CKA) (Kornblith et al., 2019), we also explored Cosine Similarity and Singular Vector Canonical Correlation Analysis (SVCCA) (Raghu et al., 2017). For brevity, we provide a brief overview of CKA below, while detailed descriptions of the remaining metrics can be found in Appendix B.
Representational Similarity using CKA.
@@
where HSIC denotes the Hilbert-Schmidt Independence Criterion (Gretton et al., 2005).
Figure 1: For each identified block of layers exhibiting redundant representations (red), we remove the internal layers of the block by zeroing out their weights (blue). For example, if a block consists of five layers, we prune layers 2 through 4, retaining only the first and last layers to reduce representation redundancy while maintaining model integrity. More details on pruning can be found in App. C.
@@
Pruning TSFMs Based on Representational Similarity.
Large TSFMs typically learn redundant representations, which often manifest as block-like structures in heatmaps depicting pairwise similarity between layer activations (Figure 1). We can leverage this redundancy to prune TSFMs, enhancing inference speed while preserving accuracy. We build on prior work (Nguyen et al., 2021) and propose a simple layer pruning strategy, which we call Block-wise Pruning, outlined in Algorithm 1. To preserve the structural integrity of each block, we retain the first and last layers of each block while zeroing out the weights of the intermediate layers. The skip connections within transformer blocks ensure that signals and gradients continue to flow through the rest of the network.
@@
  for each block $b_{i}$ in $\mathcal{B}$ do
     Let $b_{i}$ consist of layers $l_{s}$ to $l_{e}$Block edges at $l_{s}$ and $l_{e}$}
     for layer index $j=s+1$ to $e-1$ do
        Zero out the weights of layer $l_{j}$ in model $\mathcal{M}$
     end for
  end for
  return $\mathcal{M}^{\prime}$
@@
To demonstrate the effectiveness of our proposed pruning strategy, we explore two pruning configurations, one in which we prune all redundant blocks, and the other where prune only a single block. We compare the performance of these pruned models to the original, unpruned TSFMs using standard task-specific accuracy metrics (Mean Squared Error and Mean Absolute Error) and efficiency metrics (inference time in milliseconds and theoretical model size in megabytes). We evaluate these models on widely used imputation (Zhou et al., 2021) and forecasting (Ansari et al., 2024) benchmarks in both zero-shot settings and after linear probing (Goswami et al., 2024).
@@
Figure 2: Overview of linear probing, concept localization, and steering. During linear probing we train linear classifiers $f_{ij}(\mathbf{h}_{i}^{(j)},\theta_{i}^{(j)})$ to classify time series $\mathbf{x}$ into constant $c$ and sinusoid $s$ classes, using hidden representations $\mathbf{h}_{i}^{(j)}$ at the $i$-th layer and $j$-th token. We localize concepts using Fisher’s Linear Discriminant Ratio (LDR) between the classes at each layer and token. The concept steering vector at the $i$-th layer is defined as the difference between the median activation matrices of the sinusoid and constant time series classes, $\mathbf{M}_{i_{s}}-\mathbf{M}_{i_{c}}$. Vectors of all layers are stacked into a steering matrix $\mathbf{S}$ to steer model predictions towards desired concepts by updating the embeddings as $\mathbf{h}_{i}\leftarrow\mathbf{h}_{i}+\lambda\mathbf{S}_{i}$, where $\lambda$ is a scalar that controls the strength of the intervention.
Refer to caption Refer to caption Refer to caption
@@
Figure 3: Examples of synthetic data generated for concept localization, and concept steering experiments include constant signals with varying trend (left), sinusoidal signals with varying frequency (middle), and compositions of constant signals with varying trends and sinusoidal signals, resulting in sinusoidal signals with varying trends (right).
3.2 Identifying and Localizing Time Series Concepts
Through our experiments, we aim to answer the following research questions: (RQ4) Do TSFMs represent concepts associated with specific data-generating functions distinctly in the latent space? (RQ5) Are these learned concepts localized to specific layers and tokens within TSFMs?
To systematically explore the ability of TSFMs to understand intuitive time series concepts, we randomly generate a large number of synthetic univariate time series. Each randomly generated time series belong to one of two pattern classes: constant or sinusoidal. Constant patterns, represented by $y(t)=mt+b$, capture long-term non-periodic trends. Sinusoidal patterns, modeled as $y(t)=a\sin\left(\frac{2\pi t}{f}\right)$, represent periodic processes. By controlling the parameters $m$, $b$, $a$, and $f$, we can systematically generate time series with varying slope, intercept, amplitude, and periodicity, respectively. Despite their simplicity, these data generation mechanisms capture a wide range of real-world time series patterns. Example series are shown in Fig. 3. For a detailed description of the data generation process, please refer to Appendix A.
@@ 3.2 Identifying and Localizing Time Series Concepts
We build on the investigation approach outlined in (Marks & Tegmark, 2023). We say that a feature is linearly represented in a foundation model $\mathcal{M}$ if it is represented as a direction in its latent space. If this feature is linearly represented in $\mathcal{M}$, we also want to identify which layer $l$ in $\mathcal{M}$ learns this concept in the most discriminant way. As a concrete example, consider that we want to check whether $\mathcal{M}$ can distinguish between constant and sinusoidal patterns.
To determine whether the feature (sinusoidal vs. constant time series) is linearly represented, we leverage the aforementioned synthetic dataset which comprises of multiple sinusoids and constant time series randomly sampled using our data generating function. Using this dataset, we extract the residual stream of each transformer block after the feed-forward layer. Let $\mathbf{h}_{i}^{(j)}\in\mathbb{R}^{n\times D}$ denote the hidden representation of a time series $\mathbf{x}$ at $i$-th layer and $j$-th token of $\mathcal{M}$, where $D$ is the dimensionality of the hidden layer. Linear probing involves training separate linear models for each layer and token to classify time series $\mathbf{x}$ as a constant or sinusoid pattern. Classifiers $f_{ij}(\mathbf{h}_{i}^{(j)},\theta_{i}^{(j)})$ are trained on the hidden representation $\mathbf{h}_{i}^{(j)}$ at each $i$-th layer and each $j$-th token to update the parameters $\theta_{i}^{j}$. Additionally, we perform probing on representations averaged along the token dimension for each $i$-th layer. The linear probes are trained to optimize the Fisher Criterion, a function that aims to maximize the distance between class means while minimizing within-class variance:
@@
Here, $\mu_{s}$ and $\mu_{c}$ correspond to the mean computed across the $n$ dimension for each class. Similarly, $\sigma^{2}_{s}$ and $\sigma^{2}_{c}$ correspond to the variance computed across the sample dimension $n$ for each class. Let $\mathbf{V}=[v_{i,j}]\in\mathbb{R}^{L\times N}$ be the matrix of LDR values, where $v_{i,j}$ represents the LDR value for the $i$-th layer and $j$-th token, with $l$ layers and $N$ tokens. The LDR output is scaled between 0 and 1 using min-max scaling to allow for consistent comparison across layers.
By visualizing the scaled LDA values as shown in Figure 2, one can identify which layers and tokens exhibit the highest degree of separation between classes, offering insights into the network’s internal representations for concept intervention techniques.
3.3 Concept-Informed Predictions via Model Steering
Through our experiments, we aim to answer the following research questions: (RQ6) Can we leverage these learned concepts to guide model output toward concept-informed predictions? For example, can we add periodicity or an upward trend to a constant time series? (RQ7) Is it possible to combine multiple steering interventions to manipulate model predictions towards complex compositions of various concepts? For instance, can we steer a model to add both trend and periodicity to a constant signal?